# Charges, complex structures, and perturbations of instantons

### v2

@@ Appendix A Conformal-GHP connections
Consider the self-dual 2-form $Z_{ab}\equiv Z^{3}_{ab}$ given in (2.1). Its principal spinors are $o_{A},\iota_{A}$. The tensor field $J^{a}{}_{b}\equiv\sqrt{2}Z_{bc}g^{ca}$ is an almost-complex structure compatible with $g_{ab}$. Note that $J^{a}{}_{b}$ is invariant under the “GHP” transformation
@@
A spinor/tensor field $\varphi$ which transforms under (A.1)-(A.2) as $\varphi\to\Omega^{2}z^{p}\varphi$ is said to have conformal weight $w$ and GHP weight $p$. These ‘weighted’ fields are to be thought of as sections of vector bundles over $M$. A linear connection $\mathcal{C}_{AA^{\prime}}$ on these bundles was constructed in [6]. If $\phi_{AB...K}$ is a totally symmetric spinor field with conformal weight $w$ and GHP weight $p$, we have:
@@
$-f_{BA^{\prime}}\phi_{AC...L}-f_{CA^{\prime}}\phi_{BA...L}-...-f_{LA^{\prime}}\phi_{BC...A},$
@@
$f_{a}=-\frac{1}{2}J^{c}{}_{b}\nabla_{c}J^{b}{}_{a},\qquad P_{a}=\frac{i}{2}(C_{a}-J^{b}{}_{a}f_{b}),$ (A.4)
and $C_{a}$ was defined in (2.5c). Below we collect a few facts needed in the main text:
@@ Appendix B Some details of calculations
^$\hat{\nabla}_{a}\hat{V}^{a}=|\hat{j}|_{\hat{g}}^{2}-\frac{1}{2}|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}+\hat{\lambda}_{3}-\frac{\hat{R}}{6}.$ (B.2)
@@
^$\hat{\nabla}_{AA^{\prime}}\hat{V}^{AA^{\prime}}={}$ (^$(\hat{\nabla}_{AA^{\prime}}\hat{\varphi}^{AC})(\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC})+\hat{\varphi}^{AC}\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC}$
@@
[…]
where we are using the convention (2.7) for the action of the curvature operator ^$\widehat{\Box}_{AB}$. Now choose $\hat{\varphi}_{AB}$ to be a unit eigenspinor $\hat{\Psi}_{AB}{}^{CD}$ with eigenvalue $\hat{\lambda}_{3}$:
@@
Using ^$\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{EA^{\prime}}=\frac{1}{2}\delta_{A}^{E}\widehat{\Box}+\hat{\epsilon}^{EF}\widehat{\Box}_{AF}$, and the action of ^$\widehat{\Box}_{AF}$ given in (2.7), we have:
@@
The term in the second line can be rewritten using [39, Eq. (3.3.55)], after which we get
@@
[…]
@@
[…]

### v3

@@ Appendix A Conformal-GHP connections
Consider the self-dual 2-form $Z_{ab}\equiv Z^{3}_{ab}$ given in (2.1). Its principal spinors are $o_{A},\iota_{A}$. The tensor field $J^{a}{}_{b}\equiv\sqrt{2}Z_{bc}g^{ca}$ is an almost-complex structure compatible with $g_{ab}$. Note that $J^{a}{}_{b}$ is invariant under the ``GHP'' transformation
@@
A spinor/tensor field $\varphi$ which transforms under (A.1)-(A.2) as $\varphi\to\Omega^{2}z^{p}\varphi$ is said to have conformal weight $w$ and GHP weight $p$. These `weighted' fields are to be thought of as sections of vector bundles over $M$. A linear connection $\mathcal{C}_{AA^{\prime}}$ on these bundles was constructed in [7]. If $\phi_{AB...K}$ is a totally symmetric spinor field with conformal weight $w$ and GHP weight $p$, we have:
@@
$-f_{BA^{\prime}}\phi_{AC...L}-f_{CA^{\prime}}\phi_{BA...L}-.-f_{LA^{\prime}}\phi_{BC...A},$
@@
$f_{a}=-\frac{1}{2}J^{c}{}_{b}\nabla_{c}J^{b}{}_{a},\qquad P_{a}=\frac{i}{2}(\Gamma_{a}{}^{2}{}_{1}-J^{b}{}_{a}f_{b}),$ (A.4)
and $\Gamma_{a}{}^{2}{}_{1}$ was defined in (2.5c). Below we collect a few facts needed in the main text:
@@ Appendix B Some details of calculations
$\hat{\nabla}_{a}\hat{V}^{a}=|\hat{j}|_{\hat{g}}^{2}-\frac{1}{2}|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}+\hat{\lambda}_{3}-\frac{\hat{R}}{6}.$ (B.2)
@@
$\hat{\nabla}_{AA^{\prime}}\hat{V}^{AA^{\prime}}={}$ $(\hat{\nabla}_{AA^{\prime}}\hat{\varphi}^{AC})(\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC})+\hat{\varphi}^{AC}\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC}$
@@
$={}$ $|\hat{j}|_{\hat{g}}^{2}+\frac{1}{2}\hat{\varphi}^{AB}\widehat{\Box}\hat{\varphi}_{AB}-\frac{\hat{R}}{6}|\hat{\varphi}|_{\hat{g}}^{2}+\hat{\Psi}_{ABCD}\hat{\varphi}^{AB}\hat{\varphi}^{CD},$
where we are using the convention (2.7) for the action of the curvature operator $\widehat{\Box}_{AB}$. Now choose $\hat{\varphi}_{AB}$ to be a unit eigenspinor $\hat{\Psi}_{AB}{}^{CD}$ with eigenvalue $\hat{\lambda}_{3}$:
@@
Using $\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{EA^{\prime}}=\frac{1}{2}\delta_{A}^{E}\widehat{\Box}+\hat{\epsilon}^{EF}\widehat{\Box}_{AF}$, and the action of $\widehat{\Box}_{AF}$ given in (2.7), we have:
@@
The term in the second line can be rewritten using [50, Eq. (3.3.55)], after which we get
@@
$={}$ $\frac{1}{\hat{\lambda}_{3}}(3\hat{\lambda}_{3}^{2}-(2\hat{\lambda}_{1}^{2}+2\hat{\lambda}_{2}^{2}-\hat{\lambda}_{3}^{2}))$
$={}$ $\frac{1}{\hat{\lambda}_{3}}(3\hat{\lambda}_{3}^{2}-(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2})$
@@
$={}$ $\frac{2}{\hat{\lambda}_{3}}\hat{\Psi}_{ABCD}(\hat{\varphi}^{CD}\widehat{\Box}\hat{\varphi}^{AB}+\hat{\nabla}^{e}\hat{\varphi}^{AB}\hat{\nabla}_{e}\hat{\varphi}^{CD})$
$={}$ $2\hat{\varphi}_{AB}\widehat{\Box}\hat{\varphi}^{AB}+\frac{2}{\hat{\lambda}_{3}}\hat{\Psi}_{ABCD}\hat{\nabla}^{e}\hat{\varphi}^{AB}\hat{\nabla}_{e}\hat{\varphi}^{CD}$
$={}$ $-2|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}+\frac{2}{\hat{\lambda}_{3}}\hat{\Psi}_{ABCD}\hat{\nabla}^{e}\hat{\varphi}^{AB}\hat{\nabla}_{e}\hat{\varphi}^{CD}.$